Winkin, Blinkin, and Nod are equal shareholders in SleepEZ, an S corporation. In the conditions listed below, how much income should each report from SleepEZ for 2018 under both the daily allocation and the specific identification allocation method? Refer to the following table for the timing of SleepEZ’s income. Period Income January 1 through April 1 (91 days) $ 188,000 April 2 through December 31 (274 days) 361,000 January 1 through December 31, 2018 (365 days) $ 549,000 (Do not round intermediate calculations. Round your final answers to the nearest whole dollar amount.) a. There are no sales of SleepEZ stock during the year. b. On April 1, 2018, Blinkin sells his shares to Nod. c. On April 1, 2018, Winkin and Nod each sell their shares to Blinkin.
Answer #1
since there is no change of ownership both methods
will yield same result
so 549,000*1/3
183000
Daily allocation Specific identification
method method
Winkin 183,000 183,000
Blinkin 183,000 183,000
Nod 183,000 183,000
b) Daily allocation method
Winkin (549,000*365/365*1/3)=183,000
Blinkin (549000*91/365*1/3)= 45625
Nod      (549000*365/365*1/3 )+(549000*274/365*1/3)= 320375
Specific identification method
Winkin (188000*1/3)+(361000*1/3)= 182000
Blinkin (188000*1/3)+(361000*0/3)= 62667
Nod (188000*1/3)+(361000*2/3)= 303333
Daily allocation Specific identification
method method
Winkin 183,000

182,000

Blinkin 45,625 62,667
Nod 320,375 303,333
c) Daily allocation method
Winkin (549000*91/365*1/3)= 45625
Blinkin(549000*91/365*1/3)+(549000*274/365*3/3)= 457751
Nod (549000*91/365*1/3) 45625
Specific identification method
Winkin (188000*1/3)= 62666.66
Blinkin (188000*1/3)+(361000*3/3)=

423666.66

Nod (188000*1/3)= 62666.66
Daily allocation Specific identification
method method
Winkin 45,625 62,667
Blinkin 457,751 423,667
Nod 45,625 62,667
(due to rounding there is difference of 1 , please check where you need to
adjust the differecnce, I have provided the answer with full explanation